OpenStudy (anonymous):

Write the product in standard form. ( 7 + 7i)( 6 + 7i) I need help with steps

OpenStudy (bahrom7893):

just multiply it out.. 7*6 + 7*7i ...

OpenStudy (anonymous):

I don't factor?

OpenStudy (bahrom7893):

no, expand

OpenStudy (anonymous):

42+49i?

OpenStudy (bahrom7893):

well expand the whole thing...I just started it off for you.

OpenStudy (skullpatrol):

recall the FOIL method for multiplying binomials?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

( 7 + 7i)( 6 + 7i) 42 + 49i + 42i + 49i*i but i*i =1 42 + 91i +49 91 + 91i

OpenStudy (skullpatrol):

$$\Huge i*i=-1$$

OpenStudy (anonymous):

@skullpatrol : you r right so the answer will be 91i - 7

OpenStudy (skullpatrol):

@sam801 please don't give away the answer, instead let the student learn.

OpenStudy (anonymous):

Ok sorry

OpenStudy (anonymous):

i am new to this website

OpenStudy (anonymous):

so i am supposed to give only hints?

OpenStudy (anonymous):

how does it = 91-7 if you have 91+91i

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

@nieskaa : check the second line, the answer lies there

OpenStudy (anonymous):

Oh okay I got it

OpenStudy (anonymous):

@skullpatrol: thanks for the information, i just started to participate yesterday.

OpenStudy (skullpatrol):

Giving hints is a great way to find out how much the student understands.

OpenStudy (anonymous):

Sure, i will from now